OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

-1

OpenStudy (ashley_f97):

f(x) = x + 4, find f(3) f(3) = (3) + 4 =3+4 =What's 3+4?